Suppose that you decide to borrow ​$17,000 for a new car. You can select one of the following​ loans, each requiring regular monthly payments.
Installment Loan​ A: three-year loan at 5.95%
Installment Loan​ B: five-year loan at 6.46%
a. Find the monthly payments and the total interest for Loan A.
The monthly payment for Loan A is ​$
​(Do not round until the final answer. Then round to the nearest cent as​ needed.)
The total interest for Loan A is ​$
​(Round to the nearest cent as​ needed.)
b. Find the monthly payments and the total interest for Loan B.
The monthly payment for Loan B is ​$
​(Do not round until the final answer. Then round to the nearest cent as​ needed.)
The total interest for Loan B is ​$
​(Round to the nearest cent as​ needed.)
c. Compare the monthly payments and the total interest for the two loans.
Determine which loan is more economical. Choose the correct answer below.
A.
The​ five-year loan at 6.46% is more economical.
B.
The​ three-year loan at 5.95% is more economical.
The buyer will save approximately ​$ in interest.   
​(Round to the nearest cent as​ needed.)

Answer by Theo(13342)

the monthly payments on the 3 year loan are 516.787888

the monthly payments on the 5 year loan are 332.3061016

total interest on the 3 year loan is 36 * 5176.787888 - 17000 = 1604.363967.

total interest on the 5 year loan is 60 * 332.3061016 - 17000 = 2938.366094.

total interest on the 5 year loan is 2938.366094 - 1604.363967 = 1334.002128 more than on the 3 year loan.

the 3 year loan is far superior in terms of total interest owed.

i used the texas instruments business analyst 2 (TI BA II Plus).

for the 3 year loan, my inputs were.

present value = 17000
future value = 0
number of time periods = 3 years * 12 = 36 months.
interest rate per time period = 5.95% per year / 12 = .495833333% per month.
payments are made at the end of time period.

for the 5 year loan, my inputs were.

present value = 17000
future value = 0
number of time periods = 5 years * 12 = 60 months.
interest rate per time period = 6.43% per year / 12 = .535833333% per month.
payments are made at the end of each time period.

please note that, if you enter the present value as positive, the payments will show up as negative, and if you enter the present value as negative, the payments will show up as positive.

it's a cash flow convention.

i showed both as positive, since that's a human convention.